Circular No. 5980

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

GRO J0422+32 J. M. Bonnet-Bidaud, Centre d’Etudes Nucleaires de Saclay; and M.

Mouchet, Observatoire de Meudon, report: “Optical spectra of this object obtained at the Canada-France-Hawaii 3.6-m telescope on Feb. 4.28–4.36 UT confirm that the source has returned to a quiescent phase following successive optical outburts (IAUC 5974, 5967, 5929). The fully-calibrated spectrograms (range 320–1000 nm, resolution 1 nm), obtained under subarcsecond seeing conditions, show that the strong, broad Balmer lines are now observed in emission, typical of the quiescent phase of x-ray novae. Emission due to He II 468.6-nm (IAUC 5907) has disappeared. Measurements of the Hα (656.2 nm) and Hβ (486.1 nm) lines give intensities (10^{−16} erg cm^{−2} s^{−1}), equivalent widths (nm), and FWHM (km/s) of 7.64 ± 0.17, 7.7 ± 0.2, 1600 ± 50 and 3.5 ± 1.0, 4.1 ± 1.2, 1780 ± 400, respectively. The Hα line exhibits a variable double-peaked shape, similar to that observed for V404 Cyg and A0620–00, with a width that seems to indicate a rather inclined system. Preliminary analysis of the spectra show that the absorption lines from the secondary are only barely detectable, indicating a still-substantial contribution of the accretion disc.”

NOVA CASSIOPEIAE 1993 Further visual magnitude estimates (cf. IAUC 5978): Apr. 21.07 UT,

13.5 (L. Szentaskó, Veresegyház, Hungary; correction to observation of same date on IAUC 5978, which was actually for the peculiar variable in Sgr noted on IAUC 5976, and not for N Cas 1993); 23.06, 13.7 (T. Vanmunster, Landen, Belgium); 23.90, [13.3 (Vanmunster).

LUMINOUS RED VARIABLE IN SAGITTARIUS W. Wenzel, Sonneberg Observatory, reports that this object (cf. IAUC

5949) is not visible on 561 Sonneberg sky patrol plates taken during 1926– 1983 (limiting photographic magnitude 11–13.5).

COMET SHOEMAKER-LEVY (1994d) Total visual magnitude estimates: Apr. 7.83 UT, 15.0 (H. Mikuž,

Ljubljana, Slovenia, 0.20-m Baker-Schmidt camera + V filter + CCD); 11.21, 13.3 (C. S. Morris, Pine Mountain Club, CA, 0.26-m reflector).

1994 April 27 Daniel W. E. Green